# Simulating jet launching in binary neutron star mergers with no artificial atmosphere limitations

### v1

@@ Abstract
We present general relativistic magnetohydrodynamic simulations of binary neutron star (BNS) mergers, exploring the conditions to launch relativistic outflows compatible with gamma-ray burst (GRB) jets. Employing an unprecedentedly low numerical density floor decreasing as the sixth power of radial distance, we obtain for the first time an evolution where such a floor has no effect on any outflow produced during and after merger. In models forming an accreting black hole (BH)  $\simeq\,25$ ms after merger, incipient jets are launched with terminal Lorentz factors and Poynting-flux luminosities compatible with GRBs. The rather high jet velocities reached in a few tens of ms, when compared to previous simulations adopting a much higher and uniform density floor, show the importance of removing any influence of the latter as the jet evolves toward increasingly large time and spatial scales. In absence of collapse, we find the emergence of much heavier and slower polar outflows, incapable of producing a GRB. We thus favor a BH origin for GRB jets.
@@ I Introduction
The coincident detection of gravitational waves (GWs) and electromagnetic (EM) radiation from the binary neutron star (BNS) merger event GW170817 [1, 2, 3, 4, 5, 6] conclusively demonstrated that BNS mergers can power short gamma-ray burst (GRB) jets. However, modeling jet formation in merger simulations to reliably connect theoretical models to observed GRB-related EM signals remains a challenging endeavor.
Previous numerical studies employing general relativistic magnetohydrodynamic (GRMHD) simulations have shown that incipient jets (i.e. collimated, mildly relativistic, and magnetically dominated outflows) can be effectively launched by black hole-disk (BH-disk) systems formed in merger aftermaths, displaying promising terminal Lorentz factors and Poynting-flux luminosities (e.g., [7, 8, 9, 10]). Yet, the formation of an observationally consistent, mature GRB jet is not guaranteed; it critically depends on the jet’s interaction with the surrounding environment and its further evolution over larger spatial and temporal scales. In exploring the transition from promising BH-disk outflows to fully developed, high Lorentz factor jets, the presence of an excessively high atmospheric density floor employed in simulations represents a major limitation. In this work, we overcome this limitation by conducting GRMHD BNS merger simulations with an unprecedentedly low atmospheric density floor, scaling with radial distance as $\propto\!r^{-6}$. This enables a self-consistent evolution where all relevant outflows—from dynamical ejecta to incipient jets—always remain well above the atmospheric threshold.
In scenarios resulting in BH-disk formation, we obtain incipient jets exhibiting properties compatible with GRB phenomenology, and capable of propagating through the baryon-loaded polar outflows expelled prior to collapse by the highly magnetized massive neutron star (NS) remnant (e.g., [11]). [^1] For sufficiently long NS remnant lifetimes, such pre-collapse outflows can have major impact on jet propagation (e.g., [12] and refs. therein), and be potentially responsible for distinctive, relatively unexplored EM signals preceding the main GRB emission (possibly connected to observed GRB precursors [13, 14, 15]). Even with the relatively short evolution considered (up to 100 ms post-merger), the advantages of employing a low-density atmosphere become evident, particularly when comparing to previous works. Our results reinforce the expectation that, as the evolution proceeds on larger scales, a sufficiently low density atmosphere represents a necessary ingredient for a reliable description of emerging jets and their connection to GRB-related signals.
Additionally, we explore a scenario in which no collapse occurs, thereby probing the viability of the ‘magnetar’ GRB model. Consistent with [11] and a number of subsequent investigations [16, 17, 18, 19, 20, 21, 22], our simulations confirm the natural emergence of collimated outflows. However, such outflows appear incompatible with an incipient GRB jet, supporting the prevalent (but still debated) viewpoint that typical short GRBs likely require BH central engines. Notably, employing a much lower and more realistic atmospheric floor has negligible impact for outflows powerd by massive NS remnants. We conclude that removing the limitations associated with the artificial atmosphere in BNS merger simulations further favors a BH origin for GRB jets produced in this context.
@@ II Physical Models and Numerical Setup
We consider an equal-mass BNS system with individual gravitational masses of $1.3624\,M_{\odot}$, corresponding to a chirp mass of $1.186\,M_{\odot}$ (the same as the GW170817 event [23]). The initial data are generated with the LORENE library, with coordinate separation of 45 km and adopting a piecewise-polytropic approximation of the APR4 equation of state (EOS), as implemented in [24]. For the evolution, we consider a hybrid EOS obtained by adding a thermal component to the ‘cold’ piecewise-polytropic part, with the resulting pressure dependence on rest-mass density and specific internal energy given as
@@
For rest-mass densities above $10^{10}$ g/cm3, we adopt the widely used adiabatic index $\gamma_{\rm th}\!=\!1.8$ (e.g. [24] and refs. therein), while for $\rho\!<\!10^{10}$ g/cm3 we set $\gamma_{\rm th}\!=\!4/3$. The latter choice is motivated by the fact that realistic temperature-dependent EOS have an effective ideal-gas adiabatic index of 4/3 at those low densities [25] (see Supplemental material).
@@
Dipolar magnetic fields are initialized onto the two NSs at two orbits prior to merger, using the vector potential prescription of [16] with fields also extending outside the NS surface. The field strength at the NS poles is set as $B_{\rm pole}\!\simeq\!10^{15}$ G, corresponding to a total magnetic energy of $E_{\rm mag}\simeq 3\times 10^{49}$ erg at merger time. While these pre-merger magnetic fields are unrealistically strong, they still represent a small perturbation. Such field strengths are chosen to achieve the high magnetization levels expected in the post-merger phase, despite the fact that the corresponding amplification mechanisms are not fully resolved. This is a common approach in studies presenting jet-forming BNS merger simulations (e.g. [8, 10]), which are all limited by insufficient resolution.
@@
Our simulations are performed using the GRMHD code Spritz [26, 27], employing the generalised Lorenz gauge to evolve the vector potential [28], PPM reconstruction along with HLLE Riemann solver, RK4 method for time-stepping, and the publicly available RePrimAnd library for primitive variable recovery [29, 30]. Spacetime evolution follows the BSSN formulation as implemented in the MacLachlan module of the Einstein Toolkit [31]. Our Cartesian grid setup includes twelve refinement levels, with the two finest levels tracking each NS during the inspiral phase. At merger, we transition to fixed mesh refinement, and upon collapse to a BH we cover the latter with a cubic shaped extra level with $\simeq\!9.5$ km side. The computational domain extends up to 51,500 km in positive and negative $x,y,z$ directions.
@@
A key novelty compared to previous work concerns the setting of the artificial atmosphere, i.e. the minimum rest-mass density $\rho_{\mathrm{atm}}$ allowed during evolution. We recall that in Spritz, whenever the evolution would lead to $\rho\!<\!\rho_{\mathrm{atm}}$, density is reset to $\rho_{\mathrm{atm}}$, pressure is reset according to the cold part of the EOS, fluid velocity is set to zero, and magnetic fields remain unchanged. Here, we set $\rho_{\mathrm{atm}}\!=\!\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3 up to a radial distance of $r^{*}\!=\!74$ km, while at larger distances we employ the power-law profile $\rho_{\mathrm{atm}}\!=\!\rho^{*}\,(r/r^{*})^{-6}$. With this setting, no atmospheric floor reset occurs at any time above $\approx\!90$ km distance, leaving all outflows, from dynamical ejecta to the incipient jet, completely unaffected. [^2] Atmospheric densities decreasing with distance were already employed, e.g., in [32] There, $\rho_{\mathrm{atm}}\!\propto\!r^{-3}$ down to 0.166 g/cm3 and then the floor remains constant, resulting in $\rho_{\mathrm{atm}}$ higher than ours above $\sim\!700$ km. At 2000 km distance, it is already $\sim\!1000$ times higher (see comparison in Figure 1, top panel).
@@
We present three different simulations. In case A, in order to explore the scenario in which a BH-disk system powers an incipient jet, we manually induce the collapse of the NS remnant at $\approx\!25$ ms after merger, employing the procedure introduced in [33]. In case B, we consider the non-collapsing scenario. In both cases, evolved up to 75 ms after merger, the finest grid spacing is set to $dx\!=\!177$ m ($88.5$ m in the extra level covering the BH in case A). Case C is a lower resolution version of case A, with finest spacing $dx\!=\!200$ m ($100$ m around the BH), evolved up to 100 ms after merger.
Figure 1: Rest-mass density (top) and radial velocity (bottom) along the positive $x$ and $z$ axes at 20 ms after merger, for case A. Rest-mass density floor profiles in the top panel are shown as a black dashed line for our simulations and a blue dotted line for [32] (see text).
@@ III Results
In all cases, the merger produces outflows of unbound material (i.e. dynamical ejecta) in the form of equatorial tidal tails and the more isotropic shock-driven ejecta. The latter, which may affect at later times the propagation and observational signatures of an emerging jet, are characterized by a very fast low-density tail reaching velocities of $\approx\!0.9\,c$ (see Figure 1). Notably, following dynamical ejecta up to their natural front is made possible by the very low and decaying atmospheric floor.
@@
The merger results in a differentially rotating massive NS. At 20 ms post-merger, a helical magnetic tower has formed, extending to over 500 km along the NS spin axis, where magnetic field strength is $\sim\!10^{14}$ G. This tower originates polar outflows with velocities up to $\simeq\!0.2\,c$. Starting from $\simeq\!25$ ms after merger, the collapse to a BH in case A (and C) hampers the further development of such outflows and leads to a very different evolution with respect to case B, in which no collapse occurs.
@@
In cases A and C, the newly formed BH has a mass of $M_{\rm BH}\!\simeq\!2.3\,M_{\odot}$, a dimensionless spin of $\chi\!\equiv\!J/M^{2}\!\simeq\!0.61$, and is surrounded by an accretion disk of mass $M_{\rm disk}\!\simeq\!0.2\,M_{\odot}$. By the end of the simulation, $M_{\rm BH}\!\simeq\!2.4\,M_{\odot}$, $\chi\!\simeq\!0.65$, $M_{\rm disk}\!\simeq\!0.08\,M_{\odot}$, and the accretion rate has stabilized to $\dot{M}\!\simeq\!0.5\,M_{\odot}$/s. Right after collapse, as matter close to the BH is accreted, polar funnels are rapidly escavated along the spin axis. However, magnetic flux within the helical magnetic tower is in good part preserved and, in a few ms, radial velocities in the polar regions and above $\sim\!100$ km distance start to reverse. This leads to a new bipolar highly collimated outflow developing along the BH spin axis, which then matures into an incipient jet around 40 ms after merger.
@@
In Figure 2, we show the magnetic field configuration along with rest-mass density distribution for case A, at about 25 ms and 50 ms after BH formation. At both times, the average magnetic field strength around 500 km distance along the spin axis is $\sim\!10^{14}$ G, while the (mostly toroidal) field close to the BH and in the accretion disk has a strength in the range $10^{15}-10^{16}$ G. The inset of Fig. 2 gives a meridional view of the magnetization level, in terms of the magnetic energy density to rest-mass density ratio $b^{2}/(2\rho)$. For $b^{2}/(2\rho)\!\gg\!1$, this quantity also gives an estimate of the terminal lorentz factor $\Gamma_{\infty}$. Close to the BH, values $\gtrsim 10^{3}$ are achieved, which is consistent with an outflow capable of evolving into a GRB jet. Similar $\Gamma_{\infty}$ are achieved in case C.
@@
Figure 3 (top-panel) shows the time evolution of the total magnetic energy for the three cases. The initial exponential growth associated with the Kelvin-Helmholtz (KH) instability is followed by further amplification sustained by the massive NS differential rotation, via magnetic winding and the magnetorotational instability (MRI) (e.g., [34] and refs. therein). We caution, however, that our resolution is insufficient to fully capture the KH instability as well as the MRI (and the possible $\alpha\Omega$ dynamo driven by it [19]). While in case B magnetic energy keeps growing up to almost $10^{52}$ erg and then starts to slowly decrease, in cases A and C the collapse leads to a sudden drop by a factor of $\approx\!60$, after which a slowly decreasing trend leads to a final value of $\simeq\!6\times 10^{49}$ erg.
@@
Figure 2: 3D rendering of magnetic field-line structure and rest-mass density (red-to-grey colors) for case A, at about $50$ ms (left) and 75 ms (right) after merger. The inset shows a meridional view of the magnetization $b^{2}/(2\rho)$ at the jet base (vertical scale of up to ±$\pm 250$ km).
@@
Figure 3: Evolution of total magnetic energy (top) and Poynting-flux luminosity across the $r\!=\!295$ km spherical surface (bottom), for all cases. The vertical yellow line marks the merger time, whereas vertical black lines mark the collapse time for case A and C, respectively.
@@
Bottom panel of Figure 3 presents the evolution of the Poynting-flux luminosity $L_{\rm EM}$ across the spherical surface $r\!=\!295$ km. Here, the first few peaks within 8 ms after merger relate to the dynamical ejecta crossing the surface. Then, the developing magnetic tower starts to drive a magnetized outflow, gradually increasing $L_{\rm EM}$. For case B, it reaches $\sim\!10^{51}$ erg/s around 25 ms after merger, and this level is maintined until the end. In case A (and C), the collapse causes a temporary decrease until $\approx\!40$ ms after merger, when the contribution of the incipient jet emerges. In the following evolution, $L_{\rm EM}$ remains rather stationary, with a final value of $\simeq 3\times 10^{50}$ erg/s that is of the order of the power expected from the Blandford-Znajek mechanism [35, 36],
@@
Focusing now on the incipient jet of case A, we find a rapid acceleration up to $\sim\!1000$ km distance, where a maximum radial velocity of $v^{r}\!\simeq\!0.77\,c$ ($0.52\,c$) is reached on the north (south) side about $40$ ms post-collapse. Evolving further, the location of the maximum $v^{r}$ stalls around the same distance, while its value lowers to $\simeq\!0.6\,c$ ($0.5\,c$). This stalling reflects the fact that the jet is still making its way across the outer part of the pre-collapse polar outflow, as shown in Figure 4. There, we also notice a temporary switch off in jet power along the southern polar region (see also the bending of the helical magnetic field structure in Figure 2). This is due to an earlier instability at the jet-environment interface, triggering a transient baryon loading episode in the funnel.
Figure 4: Meridional view of rest-mass density (top) and radial velocity (bottom) at 75 ms after merger for cases A, B, C, and 100 ms for case C (left to right).
@@
Case C shows a jet evolution similar to case A, with maximum $v^{r}$ reaching $\simeq\!0.69\,c$ before stalling. In terms of north and south differences, it appears more symmetric (Figure 4). Continuing the evolution from 75 to 100 ms after merger, the maximum velocity starts growing again, reaching $\simeq\!0.74\,c$, while the fastest part of the incipient jet advances further before stalling again towards the end. Overall, we observe an outflow piercing through the environment in a series of finite steps, rather than experiencing a continuous advancement and acceleration. After breaking out of the dense material constituting the pre-collapse polar outflows, it will then be able to fully accelerate, potentially reaching the high Lorentz factors required to power a GRB.
In order to compare with previous work, we consider the results of [8], where the merger of (similar but not identical) BNSs, endowed with dipolar magnetic fields and described with piecewise-polytropic EOS, ends up in a BH-disk system launching an incipient jet. In the most promising model (H4M2.8P), covering $\simeq\!35$ ms after collapse, the highest velocity achieved is $\simeq\!0.64\,c$. At a similar time post-collapse, our case A (C) reaches larger values of $\simeq\!0.77\,c$ ($0.69\,c$), despite having factor $\simeq\!1.9$ (2.2) lower resolution (except for the inner $\simeq\!9.5$ km refinement), factor $\sim\!10$ lower magnetic energy, and factor $\sim\!100$ lower Poynting-flux luminosity. Moreover, in the above model H4M2.8P, the collapse occurs only $\sim\!10$ ms after merger, at which time the incipient jet of cases A and C would have propagated (much more easily) across a 60% less massive environment. In a more recent work [10], the authors presented an even more promising model (ALF2-2.70-PU5.5×1016) having resolution, magnetic energy, and collapse time similar to model H4M2.8P. In this case, the maximum jet velocity is $\simeq\!0.83\,c$ (i.e. 8% more than our case A).
While a direct one-to-one comparison is not possible, we note that simulations in [8, 10] employ a uniform atmospheric floor density of order $\sim\!10^{5}$ g/cm3, which is orders of magnitude larger than in the present work (here, with $\rho_{\mathrm{atm}}\!\propto\!r^{-6}$, the floor is already $\sim\!10^{5}$ times lower at 500 km radius). [^3] Notice also that, in [8, 10], velocity remains unchanged when density is reset to the floor value (and not set to zero as in Spritz). This difference could explain how the cases presented here, despite the much longer collapse time and much lower resolution and magnetic energy, can reach velocities almost as high as the most promising model in [10], and higher than any model in [8].

### v2

# Jet–environment interaction after delayed collapse in binary neutron star mergers

@@ Abstract
We present general relativistic magnetohydrodynamic simulations of binary neutron star (BNS) mergers, where the collapse of the metastable massive neutron star (MNS) remnant leads to the production of an incipient jet having terminal Lorentz factor and Poynting-flux luminosity compatible with a short gamma-ray burst (GRB). We consider different MNS lifetimes of about 25 and 50 ms, long enough for massive polar outflows to emerge before black hole (BH) formation. The interaction of the following BH-driven jet with such polar outflows, responsible for shock heating and possible electromagnetic signatures, is self-consistently captured for the first time. Exploiting an unprecedentedly low numerical density floor scaling as $r^{-6}$, we explore the jet propagation up to distances of $\sim\!10^{4}$ km. Comparing the outcome of different MNS lifetimes, we find that the latter, by strongly affecting the propagation environment, plays a major role in determining the final properties of the escaping jet. Finally, we consider a non-collapsing case, where the MNS-driven outflow is found too dense and slow to be compatible with a GRB jet, thus favoring a BH central engine scenario.
@@ I Introduction
The coincident detection of gravitational waves (GWs) and electromagnetic (EM) radiation from the binary neutron star (BNS) merger event GW170817 [abbott2017gw170817, abbott2017multi, pian2017spectroscopic, abbott2017grb, mooley2018superluminal, lazzati2018late] conclusively demonstrated that BNS mergers can power short gamma-ray burst (GRB) jets. However, significant uncertainties remain regarding the physical conditions under which these relativistic jets can form and successfully produce a short GRB.
General relativistic magnetohydrodynamic (GRMHD) simulations currently represent the essential tool to self-consistently investigate these physical conditions. Previous GRMHD studies have demonstrated that black-hole (BH)-powered incipient jets with promising GRB-compatible properties at their base, including terminal Lorentz factors $\Gamma_{\infty}\!\gtrsim\!100$, can emerge in cases where the massive neutron star (MNS) resulting from the merger collapses within relatively short timescales (up to a few tens of ms; e.g., [Ruiz2016, Ruiz2021, sun2022jet, Bamber2025])  [^1] We also note the recent work of [hayashi2024jet] considering the prompt collapse scenario, finding jets with $\Gamma_{\infty}\sim 10$ that are likely insufficient to power typical short GRBs.. Nevertheless, realistic MNS remnants in nature may collapse to a BH over a much broader range of lifetimes and still allow for jet formation, as suggested by several observational constraints (e.g., [Siegel2019]). Prior to collapse, longer-lived remnants can launch massive polar outflows [Ciolfi2020, Moesta2020, combi2023jets, most2023impact, Kiuchi2024, bamber2024jetlike, aguilera2024delayed, musolino2024impact], which then act as critical obstacles for the propagation of a subsequent BH-driven jet (e.g., [Pavan2025] and refs. therein). The presence of such pre-collapse outflows may also be potentially responsible for distinctive, relatively unexplored EM signals preceding the main GRB emission (possibly connected to observed GRB precursors [Troja2010, Minaev2017, zhong2019precursors]).
In this work, we perform GRMHD simulations of BNS mergers where the MNS remnant collapses to a BH either $\simeq 25$ or $\simeq 50$ ms post-merger, after which a GRB-compatible incipient jet is produced. Such lifetimes are sufficient to produce substantial pre-collapse outflows, whose self-consistent interaction with the incoming BH-driven jet is shown here for the first time. Moreover, the two different collapse times are found responsible for major differences in the incipient jet propagation, demonstrating that the MNS lifetime is a central parameter when considering GRB production in BNS mergers.
Additionally, we explore a case where no collapse occurs, thereby probing the ‘magnetar’ GRB scenario. In line with previous works [Ciolfi2020, Moesta2020, combi2023jets, most2023impact, Kiuchi2024, bamber2024jetlike, aguilera2024delayed, musolino2024impact], we find the natural emergence of steady, collimated outflows, which remain however too dense and slow to qualify as GRB-like jets.
On a technical side, we employ an unprecedentedly low numerical density floor, avoiding any artificial effect of the latter on the relevant outflows, even at increasingly large scales. This represents a key requirement for reliably connecting numerical models to the observed GRB phenomenology.
@@ II Physical Models and Numerical Setup
We consider an equal-mass BNS system with individual gravitational masses of $1.3624\,M_{\odot}$, corresponding to a chirp mass of $1.186\,M_{\odot}$ (the same as the GW170817 event [Abbott2019]). The initial data are generated with the LORENE library, with coordinate separation of 45 km and adopting a piecewise-polytropic approximation of the APR4 equation of state (EOS), as implemented in [Endrizzi2016]. For the evolution, we consider a hybrid EOS obtained by adding a thermal component to the ‘cold’ piecewise-polytropic part, with the resulting pressure dependence on rest-mass density and specific internal energy given as
@@
For rest-mass densities above $10^{10}$ g/cm3, we adopt the widely used adiabatic index $\gamma_{\rm th}\!=\!1.8$ (e.g. [Endrizzi2016] and refs. therein), while for $\rho\!<\!10^{10}$ g/cm3 we set $\gamma_{\rm th}\!=\!4/3$. The latter choice is motivated by the fact that realistic temperature-dependent EOS have an effective ideal-gas adiabatic index of 4/3 at those low densities [Bauswein2010] (see Supplemental material).
@@
Dipolar magnetic fields are initialized onto the two NSs at two orbits prior to merger, using the vector potential prescription of [Moesta2020] with fields also extending outside the NS surface. The field strength at the NS poles is set as $B_{\rm pole}\!\simeq\!10^{15}$ G, corresponding to a total magnetic energy of $E_{\rm mag}\simeq 3\times 10^{49}$ erg at merger time. While these pre-merger magnetic fields are unrealistically strong, they still represent a small perturbation. Such field strengths are chosen to achieve the high magnetization levels expected in the post-merger phase, despite the fact that the corresponding amplification mechanisms are not fully resolved. This is a common approach in studies presenting jet-forming BNS merger simulations (e.g. [Ruiz2021, Bamber2025]), which are all limited by insufficient resolution.
@@
Our simulations are performed using the GRMHD code Spritz [Cipolletta2020, cipolletta2021], employing the generalised Lorenz gauge to evolve the vector potential [Farris2012], PPM reconstruction along with HLLE Riemann solver, RK4 method for time-stepping, and the publicly available RePrimAnd library for primitive variable recovery [Kastaun2021, Kalinani2022]. Spacetime evolution follows the BSSN formulation as implemented in the MacLachlan module of the Einstein Toolkit [etk2023]. Our Cartesian grid setup includes twelve refinement levels, with the two finest levels tracking each NS during the inspiral phase. At merger, we transition to fixed mesh refinement, and upon collapse to a BH (if any) we cover the latter with a cubic shaped extra level with $\simeq\!9.5$ km side. The computational domain extends up to 51,500 km in positive and negative $x,y,z$ directions.
@@
A key novelty compared to previous work concerns the setting of the artificial atmosphere, i.e. the minimum rest-mass density $\rho_{\mathrm{atm}}$ allowed during evolution. We recall that in Spritz, whenever the evolution would lead to $\rho\!<\!\rho_{\mathrm{atm}}$, density is reset to $\rho_{\mathrm{atm}}$, pressure is reset according to the cold part of the EOS, fluid velocity is set to zero, and magnetic fields remain unchanged. Here, we set $\rho_{\mathrm{atm}}\!=\!\rho^{*}\!=\!6.17\times 10^{4}$ g/cm3 up to a radial distance of $r^{*}\!=\!74$ km, while at larger distances we employ the power-law profile $\rho_{\mathrm{atm}}\!=\!\rho^{*}\,(r/r^{*})^{-6}$. With this setting, no atmospheric floor reset occurs at any time above $\approx\!90$ km distance, leaving all outflows, from dynamical ejecta to the incipient jet, completely unaffected. The $r^{-6}$ falloff is chosen following [Pavan2021, Pavan2023], which demonstrated that such a steep decline is necessary to guarantee a reliable large-scale evolution. A shallower prescription like, e.g., the one employed in [hayashi2024jet] ($\rho_{\mathrm{atm}}\!\propto\!r^{-3}$ down to 0.166 g/cm3 and then constant) yields atmosphere densities orders of magnitude higher than ours beyond $700$ km distance and $\sim\!1000$ times higher at 2000 km, eventually becoming a critical limitation at increasingly large scales (see Supplemental material).
@@
We present four different simulations, i.e. cases A-D. In case A, in order to explore the scenario in which a BH-disk system powers an incipient jet, we manually induce the collapse of the NS remnant at $\approx\!25$ ms after merger, employing the procedure introduced in [Ciolfi2019]. In case B, we consider the non-collapsing scenario. In both cases, the finest grid spacing is set to dx=177dx\!=\!177 m ($88.5$ m in the extra level covering the BH in case A). Case C is a lower resolution version of case A, with finest spacing dx=200dx\!=\!200 m ($100$ m around the BH). Case D is the same as case C except that the collapse is induced around 50 ms after merger. All cases are evolved up to 75 ms after merger, except for case C, reaching 160 ms after merger.
@@ III Results
In all models, the merger produces outflows of unbound material (i.e. dynamical ejecta) in the form of equatorial tidal tails and the more isotropic shock-driven ejecta. The latter, which may affect at later times the propagation and observational signatures of an emerging jet, are characterized by a very fast low-density tail reaching velocities of $\approx\!0.9\,c$. Notably, following dynamical ejecta up to their natural front is made possible by the very low and decaying atmospheric floor (see Supplemental material).
@@
The merger results in a differentially rotating MNS. At 20 ms post-merger, a helical magnetic tower has formed, extending to over 500 km along the NS spin axis, where magnetic field strength is $\sim\!10^{14}$ G. This tower originates polar outflows with velocities up to $\simeq\!0.2\,c$. The collapse to a BH, $\simeq\!25$ ms after merger in cases A, C and $\simeq\!50$ ms after merger in case D, hampers the further development of such outflows and leads to a very different evolution with respect to case B, in which no collapse occurs.
Table 1: Black hole (BH) and disk properties for cases A, C, and D, shown at 1 ms and at 25 ms after collapse.
1 ms after collapse 25 ms after collapse
Case $M_{\rm BH}$ $\chi$ $M_{\rm disk}$ $M_{\rm BH}$ $\chi$ $_{\rm disk} M˙\dot{M}$
$_{\odot}] [M][M_{\odot}$] $_{\odot}] [M][M_{\odot}$] $M_{\odot}\,{\rm s}^{-1}]$
A, C 2.46 0.62 0.17 2.52 0.64 0.11 1.0
D 2.42 0.58 0.2 2.45 0.59 0.16 0.9
The BH and disk properties of collapsing models A and D are summarized in Table 1, where we report values at 1 ms and at 25 ms after collapse. Results for case A apply also to case C. For a later collapse, less material is directly swallowed by the BH, leaving a more massive disk.
@@
Right after collapse, as matter close to the BH is accreted, polar funnels are rapidly excavated along the spin axis. However, magnetic flux within the helical magnetic tower is in good part preserved and, in a few ms, radial velocities in the polar regions and above $\sim\!100$ km distance start to reverse. This leads to a new bipolar highly collimated outflow developing along the BH spin axis, which then matures into a magnetically dominated incipient jet around 15 ms after collapse.
@@
In Figure 1, we show the magnetic field configuration along with rest-mass density distribution for case A, at about 25 and 50 ms after BH formation. At both times, the average magnetic field strength around 500 km distance along the spin axis is $\sim\!10^{14}$ G, while the (mostly toroidal) field close to the BH and in the accretion disk has a strength in the range $10^{15}-10^{16}$ G. The inset of Fig. 1 gives a meridional view of the magnetization level, in terms of the magnetic energy density to rest-mass density ratio $b^{2}/(2\rho)$. For $b^{2}/(2\rho)\!\gg\!1$, this quantity also gives an estimate of the terminal Lorentz factor $\Gamma_{\infty}$. Close to the BH, values $10^{3}$ are achieved, which is consistent with an outflow capable of evolving into a GRB jet. Similar values are obtained in case C, while $\Gamma_{\infty}$ reaches a few hundred in case D (see Supplement).
@@
Figure 2 (top-panel) shows the time evolution of the total magnetic energy for the four cases. The initial exponential growth associated with the Kelvin-Helmholtz (KH) instability is followed by further amplification sustained by the massive NS differential rotation, via magnetic winding and the magnetorotational instability (MRI) (e.g., [Ciolfi2020b] and refs. therein). We caution, however, that our resolution is insufficient to fully capture the KH instability as well as the MRI (and the possible $\alpha\Omega$ dynamo driven by it [Kiuchi2024]). While in case B magnetic energy keeps growing up to almost $10^{52}$ erg and then starts to slowly decrease, in cases A and C the collapse leads to a sudden drop by a factor of $\approx\!60$, after which a slowly decreasing trend leads to a final value of $\simeq\!6\!\times\!10^{49}$ erg. Case D indicates that a longer remnant lifetime enhances the outflow of magnetized material into the environment, yielding nearly twice the magnetic energy at the final time compared to cases A and C.
@@
Figure 1: 3D rendering of magnetic field-line structure and rest-mass density (red-to-grey colors) for case A, at about $50$ ms (left) and 75 ms (right) after merger. The inset shows a meridional view of the magnetization $b^{2}/(2\rho)$ at the jet base (vertical scale of up to $\pm 250$ km).
@@
Figure 2: Evolution of total magnetic energy (top) and Poynting-flux luminosity across the r=295r\!=\!295 km spherical surface (bottom), for all cases. The vertical light blue line marks the merger time, whereas vertical black lines mark the collapse time for case A, C, and D, respectively.
@@
Bottom panel of Figure 2 presents the evolution of the Poynting-flux luminosity $L_{\rm EM}$ across the spherical surface at r=295r\!=\!295 km. Here, the first few peaks within 8 ms after merger relate to the dynamical ejecta crossing the surface. Subsequently, the developing magnetic tower starts to drive a magnetized outflow, gradually increasing $L_{\rm EM}$. For case B, $L_{\rm EM}$ reaches $\sim\!10^{51}$ erg/s around 25 ms after merger, and this level is maintained until the end. In cases A, C and D, the collapse triggers a temporary drop until $\approx\!15$ ms after BH formation, after which the contribution of the incipient jet becomes apparent. In the following evolution, $L_{\rm EM}$ remains rather stationary for cases A and C, settling at $\simeq 3\times 10^{50}$ erg/s, while for case D it increases to $\simeq 10^{51}$ erg/s. These values are comparable to the power expected from the Blandford-Znajek mechanism [Blandford1977, Thorne1986],
@@
Figure 3 compares cases A (top) and C (bottom) at $t\!\simeq\!50$ ms after merger (i.e. $t\!-\!t_{\rm BH}\!\simeq\!25$ ms), showing unbound matter ($u_{t}\!>\!1$, according to geodesic criterion), internal energy density $\rho\epsilon$, and effective specific entropy $s_{\rm eff}\!\equiv\!\ln(P/\rho^{\gamma_{\rm eff}})$ computed from the effective adiabatic index $\gamma_{\rm eff}\!\equiv\!P/(\rho\epsilon)+1$. In both collapsing scenarios, the incipient jet launched by the BH–disk system is highly unbound, and its front eventually encounters the more dense polar outflow expelled during the MNS phase. This interaction leads to alternating phases of jet advancement and stalling and produces shock heating, clearly visible as localized enhancements in $\rho\epsilon$ and sharp entropy jumps around $10^{3}$ km distance and $\lesssim\!15^{\circ}$ from the $z$-axis. These shocks, while dissipating jet energy, may leave a detectable imprint in the following EM emission, possibly powering a GRB precursor signal. Moreover, by modifying the thermodynamic conditions of the involved material, they can influence nucleosynthesis and possibly affect the resulting kilonova emission.
Figure 4 shows unbound regions ($u_{t}\!>\!1$) for all cases at t=75t\!=\!75 ms, and also at $t\!\simeq\!160$ ms for case C. In cases A and C, we find rapid jet acceleration up to $\sim\!1500$ km at t=75t\!=\!75 ms, reaching peak radial velocities of $v^{r}\!\simeq\!0.77$ c and $0.69$ c, respectively. Comparing the above velocities, we find that higher resolution favors a more efficient jet acceleration, although the qualitative dynamics remains the same. In both cases, the jet front extends to $\sim\!3\!\times\!10^{3}$ km, but its velocity is slowed down to $\sim\!0.6$ c by the dense pre-collapse outflow. This stalling reflects the fact that the jet is still making its way across the outer part of the pre-collapse polar outflow.
In case D, delaying the collapse by $\simeq\!25$ ms is sufficient to strongly increase the level of baryon pollution due to the MNS-driven outflow, which results more massive and extended (Fig. 5). As a consequence, the incipient jet has to spend more time and energy to advance through the surroundings (e.g., compare, at the same time after collapse, case D in Fig. 4 and case C in Fig. 3; see also Supplemental material). Moreover, it retains orders of magnitude higher density (Fig. 5) and, in turn, lower terminal Lorentz factor. These findings show that the collapse time, by strongly affecting the properties of the environment through which the incipient jet has to propagate, plays a fundamental role in determining the final jet energetics and breakout timescale.
Finally, in case B, the unbound polar outflow produced by the MNS remains rather slow, with $v^{r}\lesssim\!0.3$ c. Moreover, its maximum terminal Lorentz factor is orders of magnitude lower compared to the BH-powered incipient jets of cases A,C, and D, and incompatible with a GRB (see also Fig. 4 and Supplemental material).
@@
Continuing the evolution of case C from 75 to 160 ms after merger (Fig. 4, rightmost panels), the maximum velocity starts growing again, reaching $\simeq\!0.74\,c$, while the fastest part of the incipient jet advances further before stalling again towards the end. At final time, the jet front extends beyond $\sim\!10^{4}$ km. Overall, we observe a jet piercing through the environment in a series of finite steps, rather than experiencing a continuous advancement and acceleration. Only after breaking out of the dense material constituting the pre-collapse polar outflows, it be able to fully accelerate, potentially reaching the high Lorentz factors required to power a GRB.
Figure 3: Meridional view of $u_{t}$ (left), internal energy density $\rho\epsilon$ (center), and effective specific entropy $s_{\rm eff}$ (right) at $t\!\simeq\!50$ ms after merger for case A (z>0z\!>\!0, top) and C (z<0z\!<\!0, bottom).
Figure 4: Meridional view of $u_{t}$ at 75 ms after merger for cases A (top-left; $z>0$), B (top-center; $z>0$), C (bottom-left; $z<0$), D (bottom-center; $z<0$), and at 162 ms after merger for case C on larger spatial scales (right).